1. When a current of 0.82A was passed for 5 hours through an aqueous solution of metal “Z” 2.65g of metal were deposited Determine the change on the ions of metal (A faraday = 96,500 coulomb) relativeatomic mass of Z = 52 ( 2 mks)
1
Expert's answer
2020-08-05T05:08:40-0400

I = 0.82 A

m = mass of Z = 2.65 g

F = 96500 Coulombs

t = 5 hours = 300 min = 18000 sec

M = Relative atomic mass of Z = 52

Q = I × t


Solution:

Faraday's First Law of Electrolysis:

"The mass of the substance (m) deposited or liberated at an electrode is directly proportional to the quantity of electricity or charge (Q) passed."

In the mathematical form, this law can be represented as follows:

m = (Q × M) / (F × z)

or

m = (I × t × M) / (F × z)


Thus:

2.65 g = (0.82 A × 18000 sec × 52) / (96500 × z)

(0.82 A × 18000 sec × 52) = (2.65 g × 96500 × z)

767520 = 255725 × z

z = +3

The charge on the ions of metal is +3.

Z3+ + 3e = Z0


Answer: The charge on the ions of metal is +3 (Z3+).